# algorithmic

### v2

@@ 5.2.1 Empirical Risk Minimization
The most prevalent variant for learning the reward function, already been presented in Section 2.3, is a special case of empirical risk minimization. The general approach of empirical risk minimization for reward function learning, assuming an underlying human feedback model with utilities as in (9) is to find the minimizer of
@@
where $\mathcal{D}={\{(l_{i},q_{i})\}}_{i=1}^{N}$ is the given data set of observed label and query pairs, :×$\ell:\mathbb{R}\times\mathcal{Q}\times\mathcal{C}$ is a suitable loss function with $\mathcal{Q}$ being the set of all possible label sets, and $u_{\cdot}(R)$ denoting the utility (depending on the return $R$) of the possible labels for the given label-query pair $l_{i},m(q_{i}).$ As an illustration, consider the common case of pairwise trajectory comparisons where queries are pairs of trajectories $q_{i}=\{\tau_{1}^{i},\tau_{2}^{i}\}$, and labels $l_{i}\in\{\tau_{1}^{i}\succ\tau_{2}^{i},\tau_{1}^{i}\prec\tau_{2}^{i}\}=m(q_{i})$ are the human’s preference over the two trajectories. For a given query $q_{i}$, we then obtain (2) as a special case of (12) by using the loss function:
@@
=$\mathcal{R}=\{R_{\bm{\psi}}(s,a)=\bm{\psi}^{\top}\bm{\phi}(s,a)\,|\,(s,a)\in\mathcal{S}\times\mathcal{A},\bm{\psi}\in\mathbb{R}^{d}\},$
@@
Since the assumption of a linear reward model may be too strong in practice, most recent work is based on non-linear models, especially using differentiable models, but other cases have been investigated as well. In the latter case, for instance, decision trees have been considered to learn an interpretable reward model (see Section 5.2.4). In the former case, simple multilayer perceptron (MLP) has naturally been considered, but more recent deep learning architectures are more commonly used in the recent literature. Thus, especially with partially observable domains, the reward network may be composed of a state-action encoder followed by fully connected layers. For instance, Abramson et al. (2022) combine ResNet blocks for image processing, a learnable embedding table, a multi-modal transformer, LSTMs, and MLPs. Besides, Kim et al. (2023) utilize a Transformer-based architecture (Vaswani et al., 2017), motivated by the observation that rewards are often non-Markovian.
@@
The main supervised loss to train the reward model can also be augmented with additional losses corresponding to auxiliary tasks to avoid overfitting and improve generalizability. For instance, Abramson et al. (2022) use a behavior cloning loss and add a policy head to the reward network, thereby preventing that the reward model drifts too far from its initialization from the policy. Metcalf et al. (2023) design a reward model using state-action representations trained to be temporally consistent via self-supervised learning. On a related note, the scalar preference optimization problem has been extended to a multidimensional one by Zhong et al. (2024) to represent diverse human preferences and Marta et al. (2023) for query efficiency.
@@ 5.2.3 Partial Identifiability
A crucial question when it comes to learning the reward function is whether the reward function can be identified at all. If two reward functions induce exactly the same human feedback model, the reward function is called partially identifiable or ambiguous. Skalse et al. (2023) study this topic for the Boltzmann distribution as the underlying human feedback model when demonstrations (inverse RL) or pairwise trajectory preferences are given as feedback. For demonstrations, this question has also been examined in other works (Ng & Russell, 2000; Dvijotham & Todorov, 2010; Kim et al., 2021; Cao et al., 2021a). On a related note, Ellis et al. (2024) tackle this identifiability issue by considering suitable acquisition functions (see Section 4.1.1).
@@ 5.2.4 Interpretability
The field of explainable artificial intelligence (XAI) has emerged in recent years to improve the transparency and explainability of models or even to enable them in the first place. Roughly speaking, the aim is to resort to more interpretable methods or provide explanations for both experts and non-experts, shedding light on why a certain input in a (black box) model leads to a certain result. Explanations can take different forms, as can the ways to ensure the transparency of models, and for a detailed overview, we refer to Barredo Arrieta et al. (2020). It is worth noting that the field has grown so extensively over the years that even dedicated overviews for the field of interpretable and explainable RL are by now available (Puiutta & Veith, 2020; Glanois et al., 2022; Qing et al., 2023; Milani et al., 2023).
For the branch of RLHF, the existing works are quite sparse and mostly limited to using tree models as transparent and explainable models for learning the reward function (Bewley & Lécué, 2022; Bewley et al., 2022; Kalra & Brown, 2022; Bewley et al., 2024; Kalra & Brown, 2023). Another way to realize explainability within RLHF suggested by Zhang & Kashima (2023) is to learn simultaneously the reward function and the importance of states using a weight network. Assuming that for (long) trajectories, only a few states are important for the preference outcome, their framework can be used to select samples for explainability purposes. Moreover, a perturbation analysis is suggested to evaluate explanations in a quantitative manner using the learned state importance weights.
@@ 5.2.5 Online Improvements
This is related to issues posed by the approximate nature of the reward model in general, discussed in further detail in Section 6.1, but improving reward model accuracy, in general, is not sufficient: McKinney et al. (2022) further show the interdependence of the reward model and the policy, demonstrating that reward models trained online together with a policy may not be effective when a completely new policy is trained.
Solutions to the problems of overoptimization and interdependence can take different forms: One is to update the reward model online with sufficient frequency using notably more on-policy queries (see Section 4.1.1), another is to improve the reward model, e.g., by leveraging ensembles (Coste et al., 2024) or by modifying the training procedure to place additional emphasis on challenging examples (Zheng et al., 2024), and a third, discussed in Section 6.1, is to add constraints to the policy training.
@@ 5.2.6 Learning from Multiple Feedback Types
As discussed in Section 3.5, it is often desirable to combine several feedback types. This requires extensions of the learning process to incorporate different sources of feedback. Learning from multiple feedback types can be achieved by pre-processing the feedback, assuming common latent factors, or by using the feedback types for distinct purposes.
The first approach is demonstrated by (Novoseller et al., 2023), who infer preferences from demonstrations, allowing them to treat both types of feedback equally in the learning pipeline. In the style of the second approach, Jeon et al. (2020) proposes the unified framework of reward-rational choice, which allows for interpreting many forms of human feedback as Boltzmann-rational choices and, through this common framework, enables combination and adaptive selection of feedback types. Finally, different types of feedback can be used for entirely different purposes, such as one for objective learning and another for safety-constraints (Cosner et al., 2022) or for representation learning (see Section 3.2.3).
Since multiple sources of reward information may conflict, it is important to consider how to combine them. Krasheninnikov et al. (2021) study several possible strategies of combining several reward functions in this setting, relating it to multi-task inverse RL. Note that this challenge of conflicting sources of reward relates tightly to challenges posed when receiving diverse preferences from different labelers, as discussed in Section 5.1.6.
@@ 5.2.7 Offline Reward Learning
There is a recent trend towards offline RLHF, where both the reward model and the policy are trained offline. The offline setting is also frequently considered in RLHF theory (Section 7). Early approaches in this area (Kim et al., 2023; Shin et al., 2023) first generate queries from an offline dataset of behaviors, gather human responses, train a reward model from the resulting preference data, and then leverage offline RL algorithms to derive a policy. We do not cover these works in detail, since this survey primarily focuses on the interactive and online setting (see Section 1.3). Nonetheless, the offline setting is particularly useful for evaluating novel approaches, e.g., for active query selection, using offline datasets. We refer the interested readers to Section 8.4 for a discussion of available datasets.
@@ 5.3 Evaluating Learned Reward Functions
When online rollouts are not possible, so-called off-policy evaluations, which estimate the value of the optimal policy on the basis of an available data set, may be considered. For coping with biases or large variances due to policy mismatch, approaches using importance sampling (Precup et al., 2000), regression- or classification-based methods (Paduraru, 2013; Le et al., 2019; Irpan et al., 2019), or combinations of these (Jiang & Li, 2016) have been proposed. The problem with these approaches, however, is that the traces of the explicit sources of error through policy learning or reward learning are blurred, and that they require access to the ground-truth rewards.
@@
Even if some attractive properties, above all a Lipschitz continuity in terms of the EPIC distance of two reward functions for the difference of the value functions of the induced optimal policies is shown, this distance has its shortcomings. One of these is that the canonicalization form used by EPIC distance does not encode sufficient knowledge about the underlying transition dynamics, which might lead to unreliable distances when evaluating reward functions on physically non-realizable transitions. To this end, Wulfe et al. (2022) propose the Dynamics-Aware Reward Distance (DARD), which uses a slightly different form of canonicalization but restricts the evaluation of the reward functions to transitions that are approximately physically feasible.
@@
For an evaluation by visual inspection, Jenner & Gleave (2021) propose a method for preprocessing reward functions by transforming them into simpler but equivalent reward functions for better interpretability. Related to this and the rollout method, the quality of the reward function learned can also be evaluated by a human (or expert) by examining the behavior of the agent on the target task. Besides, in the context of LLMs, datasets have been proposed and specifically designed to evaluate the (in)consistency of learned reward models with respect to semantic changes of prompts (Shen et al., 2024).
@@ 5.4 Reward Model Inputs
Besides the feedback type, another factor is the modality of the reward model input data. This usually consists of the agent’s observations and actions. Observations can range from true state to high dimensional inputs (e.g., images), while actions can range from discrete finite actions to continuous actions.
@@ 5.5 Increasing Feedback Efficiency
Maximizing feedback efficiency is vital in RLHF due to the high cost of human feedback. This section delves into methods that enhance learning from limited human feedback. We discuss methods that leverage prior offline data, methods that use (partially unlabelled) data more efficiently, and methods that aim to gather more informative data.
@@ 5.5.1 Using Prior Data
There are often large amounts of prior data available at little or no additional cost. While this data generally was generated for other tasks, many basic human preferences are the same for various tasks and can often even be extracted from completely unsupervised data such as text corpora. By leveraging this prior data, we can greatly reduce the amount of feedback necessary to learn the current task’s objective. We explore various methods, including meta- and transfer learning, leveraging foundation models, reward model initialization, preference model pretaining, and supervised representation learning.
@@
Foundation models, i.e., large models trained on large amounts of often unlabeled data, can acquire significant knowledge about basic human preferences. A language model trained to predict the next token in a text corpus, for example, may learn to complete the sentence ‘Frank was mad that his vacuum robot broke the vase,’ thereby learning that humans prefer non-destructive behavior. These learned preferences can then be leveraged in RLHF approaches. For instance, Kwon et al. (2023) propose to use LLM as a source of rewards. Du et al. (2023) is another example, where a success detector is trained using a pre-trained vision-language model (Flamingo). Their approach utilizes a data set of trajectories with binary success labels, employing a non-interactive training method.
@@
It is often beneficial to initialize the reward model with parameters from a model trained on a related task. This strategy is particularly common in language model fine-tuning, where self-supervised pretraining is a common practice. In such scenarios, it becomes logical to use these pre-trained models for initializing not just the policy but also the reward model. This methodology is adopted by Askell et al. (2021) and Ouyang et al. (2022). Specifically, Ouyang et al. (2022) use a pretrained language model for the reward model, opting for a smaller model relative to the policy to mitigate unstable learning. Notably, while they apply supervised fine-tuning to the policy before the RLHF phase, the reward model is initialized directly from the language model without any preliminary fine-tuning. This approach’s applicability extends beyond language models to other areas. A notable example is Abramson et al. (2022), who, in the control domain, begin by training a policy through contrastive self-supervised learning and behavioral cloning. They then add an MLP head to the policy for the prediction of cumulative rewards.
@@
A compact representation that captures all relevant information for human preferences while minimizing noise can greatly enhance preference learning efficiency. It may also generalize better than a representation learned end-to-end as part of the preference learning task, which may contain spurious correlations. Bobu et al. (2022) address this by proposing the learning of features through explicit human feedback using feature traces. Feature traces (see Section 3.2.3) involve human labelers explicitly teaching relevant features one by one by demonstrating behavior in which the feature monotonically increases or decreases. This method directly aligns the learned representation with human-identified features, enhancing preference learning efficiency but requiring detailed human input. However, feature traces require labelers to be able to identify and articulate relevant features, which can be challenging. Bobu et al. (2023) offer an alternative approach with their Similarity-based Implicit Representation Learning (SIRL) method. SIRL learns representations from similarity queries (see Section 3.2.3), where human labelers provide feedback on whether behaviors are similar or different concerning the features that matter to them. This method captures a broader range of human notions of similarity without needing explicit feature knowledge, thus reducing the cognitive load on human labelers. In summary, while both approaches emphasize human feedback’s centrality in representation learning, they differ in their methods of gathering this feedback. The feature traces used by Bobu et al. (2022) require specific feature knowledge, whereas SIRL used by Bobu et al. (2023) utilizes more intuitive similarity assessments, potentially offering a more user-friendly way to capture human preferences.
These diverse methods of utilizing prior data demonstrate the potential for enhancing data efficiency in RLHF, enabling more effective learning from limited human feedback.

### v3

@@ 5.2.1 Empirical Risk Minimization
The most prevalent variant for learning the reward function, already presented in Section˜2.3, is a special case of empirical risk minimization. The general approach of empirical risk minimization for reward function learning, assuming an underlying human feedback model with utilities as in (9) is to find the minimizer of
@@
where $\mathcal{D}={\{(l_{i},q_{i})\}}_{i=1}^{N}$ is the given data set of observed label and query pairs, $\ell:\mathbb{R}\times\mathcal{Q}\times\mathcal{C}$ is a suitable loss function with $\mathcal{Q}$ being the set of all possible label sets, and $u_{\cdot}(R)$ denoting the utility (depending on the return $R$) of the possible labels for the given label-query pair $l_{i},m(q_{i})$. As an illustration, consider the common case of pairwise trajectory comparisons where queries are pairs of trajectories $q_{i}=\{\tau_{1}^{i},\tau_{2}^{i}\}$, and labels $l_{i}\in\{\tau_{1}^{i}\succ\tau_{2}^{i},\tau_{1}^{i}\prec\tau_{2}^{i}\}=m(q_{i})$ are the human’s preference over the two trajectories. For a given query $q_{i}$, we then obtain (2) as a special case of (12) by using the loss function:
@@
$\mathcal{R}=\{R_{\bm{\psi}}(s,a)=\bm{\psi}^{\top}\bm{\phi}(s,a)\,|\,(s,a)\in\mathcal{S}\times\mathcal{A},\bm{\psi}\in\mathbb{R}^{d}\},$
@@
Since the assumption of a linear reward model may be too strong in practice, most recent work is based on non-linear models, especially using differentiable models, but other cases have been investigated as well. In the latter case, for instance, decision trees have been considered to learn an interpretable reward model (see Section˜5.2.4). In the former case, simple multilayer perceptron (MLP) has naturally been considered, but more recent deep learning architectures are more commonly used in the recent literature. Thus, especially with partially observable domains, the reward network may be composed of a state-action encoder followed by fully connected layers. For instance, Abramson et al. (2022) combine ResNet blocks for image processing, a learnable embedding table, a multi-modal transformer, LSTMs, and MLPs. Besides, Kim et al. (2023) use a Transformer-based architecture (Vaswani et al., 2017), motivated by the observation that rewards are often non-Markovian.
@@
The main supervised loss to train the reward model can also be augmented with additional losses corresponding to auxiliary tasks to avoid overfitting and improve generalizability. For instance, Abramson et al. (2022) use a behavior cloning loss and add a policy head to the reward network, thereby preventing that the reward model drifts too far from its initialization from the policy. Metcalf et al. (2023) design a reward model using state-action representations trained to be temporally consistent via self-supervised learning. On a related note, the scalar preference optimization problem has been extended to a multidimensional one by Zhong et al. (2024b) to represent diverse human preferences and Marta et al. (2023) for query efficiency.
@@ 5.2.3 Partial Identifiability
A crucial question when it comes to learning the reward function is whether the reward function can be identified at all. If two reward functions induce exactly the same human feedback model, the reward function is called partially identifiable or ambiguous. Skalse et al. (2023) study this topic for the Boltzmann distribution as the underlying human feedback model when demonstrations (inverse RL) or pairwise trajectory preferences are given as feedback. For demonstrations, this question has also been examined in other works (Ng & Russell, 2000; Dvijotham & Todorov, 2010; Kim et al., 2021; Cao et al., 2021a). On a related note, Ellis et al. (2024) tackle this identifiability issue by considering suitable acquisition functions (see Section˜4.1.1).
@@ 5.2.4 Interpretability
The field of explainable artificial intelligence (XAI) has emerged in recent years to improve the transparency and explainability of models or even to enable them in the first place. Roughly speaking, the aim is to resort to more interpretable methods or provide explanations for both experts and non-experts, shedding light on why a certain input in a (black box) model leads to a certain result. Explanations can take different forms, as can the ways to ensure the transparency of models, and for a detailed overview, we refer to Barredo Arrieta et al. (2020). It is worth noting that the field has grown so extensively over the years that even dedicated overviews for the field of interpretable and explainable RL are by now available (Puiutta & Veith, 2020; Qing et al., 2023; Milani et al., 2023; Glanois et al., 2024).
For the branch of RLHF, the existing works are quite sparse and mostly limited to using tree models as transparent and explainable models for learning the reward function (Bewley & Lécué, 2022; Bewley et al., 2022; Kalra & Brown, 2022; Bewley et al., 2024; Kalra & Brown, 2024). Another way to realize explainability within RLHF suggested by Zhang & Kashima (2023) is to learn simultaneously the reward function and the importance of states using a weight network. Assuming that for (long) trajectories, only a few states are important for the preference outcome, their framework can be used to select samples for explainability purposes. Moreover, a perturbation analysis is suggested to evaluate explanations in a quantitative manner using the learned state importance weights.
@@ 5.2.5 Online Improvements
This is related to issues posed by the approximate nature of the reward model in general, discussed in further detail in Section˜6.1, but improving reward model accuracy, in general, is not sufficient: McKinney et al. (2022) further show the interdependence of the reward model and the policy, demonstrating that reward models trained online together with a policy may not be effective when a completely new policy is trained.
Solutions to the problems of overoptimization and interdependence can take different forms: One is to update the reward model online with sufficient frequency using notably more on-policy queries (see Section˜4.1.1), another is to improve the reward model, e.g., by leveraging ensembles (Eisenstein et al., 2024; Coste et al., 2024), combining multiple reward models with weight averaging(Rame et al., 2024), or by modifying the training procedure to place additional emphasis on challenging examples (Zheng et al., 2024), and a third, discussed in Section˜6.1, is to add constraints to the policy training.
@@ 5.2.6 Learning from Multiple Feedback Types
As discussed in Section˜3.5, it is often desirable to combine several feedback types. This requires extensions of the learning process to incorporate different sources of feedback. Learning from multiple feedback types can be achieved by pre-processing the feedback, assuming common latent factors, or by using the feedback types for distinct purposes.
The first approach is demonstrated by (Novoseller et al., 2023), who infer preferences from demonstrations, allowing them to treat both types of feedback equally in the learning pipeline. In the style of the second approach, Jeon et al. (2020) proposes the unified framework of reward-rational choice, which allows for interpreting many forms of human feedback as Boltzmann-rational choices and, through this common framework, enables combination and adaptive selection of feedback types. Finally, different types of feedback can be used for entirely different purposes, such as one for objective learning and another for safety-constraints (Cosner et al., 2022) or for representation learning (see Section˜3.2.3).
Since multiple sources of reward information may conflict, it is important to consider how to combine them. Krasheninnikov et al. (2021) study several possible strategies of combining several reward functions in this setting, relating it to multi-task inverse RL. Note that this challenge of conflicting sources of reward relates tightly to challenges posed when receiving diverse preferences from different labelers, as discussed in Section˜5.1.6.
@@ 5.2.7 Offline Reward Learning
There is a recent trend towards offline RLHF, where both the reward model and the policy are trained offline. The offline setting is also frequently considered in RLHF theory (Section˜7). Early approaches in this area (Kim et al., 2023; Shin et al., 2023) first generate queries from an offline dataset of behaviors, gather human responses, train a reward model from the resulting preference data, and then leverage offline RL algorithms to derive a policy. We do not cover these works in detail, since this survey primarily focuses on the interactive and online setting (see Section˜1.3). Nonetheless, the offline setting is particularly useful for evaluating novel approaches, e.g., for active query selection, using offline datasets. We refer the interested readers to Section˜8.4 for a discussion of available datasets.
Note that offline reward learning does not necessarily need to be combined with offline policy learning, as the reward model can be used to train a policy online as well. This can be problematic in practice, as the distribution of experiences will generally shift during online policy learning, which can lead to poor performance of the reward model. Nonetheless, this setup is commonly used in the LLM finetuning setting, since distributional shift is less of a concern due to the large and diverse training data (often collected off-policy) and the bandit-like nature of the task (compare Section˜6.2). In this setting, it is also common to directly integrate preference and policy learning through direct policy optimization (DPO) techniques further discussed in Section˜6.3.
@@ 5.3 Evaluating Learned Reward Functions
When online rollouts are not possible, so-called off-policy evaluations, which estimate the value of the optimal policy based on an available data set, may be considered. For coping with biases or large variances due to policy mismatch, approaches using importance sampling (Precup et al., 2000), regression- or classification-based methods (Paduraru, 2013; Le et al., 2019; Irpan et al., 2019), or combinations of these (Jiang & Li, 2016) have been proposed. Off-policy evaluation is particularly well-suited for evaluating LLMs reward models. Lambert et al. (2024) propose a particularly large-scale dataset for this purpose, focusing on examples that are difficult to judge. They then use accuracy-based metrics across multiple categories to evaluate reward models. The problem with these approaches, however, is that the traces of the explicit sources of error through policy learning or reward learning are blurred, and that some require access to the ground-truth rewards.
@@
Even if some attractive properties, above all a Lipschitz continuity in terms of the EPIC distance of two reward functions for the difference of the value functions of the induced optimal policies is shown, this distance has its shortcomings. One shortcoming is that the canonicalization form used by EPIC distance does not encode sufficient knowledge about the underlying transition dynamics, which might lead to unreliable distances when evaluating reward functions on physically non-realizable transitions. To this end, Wulfe et al. (2022) propose the Dynamics-Aware Reward Distance (DARD), which uses a slightly different form of canonicalization but restricts the evaluation of the reward functions to transitions that are approximately physically feasible.
@@
For an evaluation by visual inspection, Jenner & Gleave (2021) propose a method for preprocessing reward functions by transforming them into simpler but equivalent reward functions for better interpretability. Related to this and the rollout method, the quality of the reward function learned can also be evaluated by a human (or expert) by examining the behavior of the agent on the target task. This can even extended to the reinforcement learning from AI feedback (RLAIF) setting, by using another trained model as the evaluator (Li et al., 2024). Besides, in the context of LLMs, datasets have been proposed and specifically designed to evaluate the (in-)consistency of learned reward models with respect to semantic changes of prompts (Shen et al., 2024).
@@ 5.4 Reward Model Inputs
Besides the feedback type, another factor is the modality of the reward model input data. This usually consists of the agent’s observations and actions. Observations can range from true state to high-dimensional inputs (e.g., images), while actions can range from discrete finite actions to continuous actions.
@@ 5.5 Increasing Feedback Efficiency
Maximizing feedback efficiency is vital in RLHF due to the high cost of human feedback. This section delves into methods that enhance learning from limited human feedback. We discuss methods that leverage prior offline data, methods that use (partially unlabeled) data more efficiently, and methods that aim to gather more informative data.
@@ 5.5.1 Using Prior Data
There are often large amounts of prior data available at little or no additional cost. While this data generally was generated for other tasks, many basic human preferences are the same for various tasks and can often even be extracted from completely unsupervised data such as text corpora. By leveraging this prior data, we can greatly reduce the amount of feedback necessary to learn the current task’s objective. We explore various methods, including meta- and transfer learning, leveraging foundation models, reward model initialization, preference model pretraining, and supervised representation learning.
@@
Foundation models, i.e., large models trained on large amounts of often unlabeled data, can acquire significant knowledge about basic human preferences. A language model trained to predict the next token in a text corpus, for example, may learn to complete the sentence ‘Frank was mad that his vacuum robot broke the vase’, thereby learning that humans prefer non-destructive behavior. These learned preferences can then be leveraged in RLHF approaches. For instance, Kwon et al. (2023) propose to use LLM as a source of rewards. Du et al. (2023) is another example, where a success detector is trained using a pre-trained vision-language model (Flamingo). Their approach uses a dataset of trajectories with binary success labels, employing a non-interactive training method.
@@
It is often beneficial to initialize the reward model with parameters from a model trained on a related task. This strategy is particularly common in language model fine-tuning, where self-supervised pretraining is a common practice. In such scenarios, it becomes logical to use these pre-trained models for initializing not just the policy but also the reward model. This methodology is adopted by Askell et al. (2021) and Ouyang et al. (2022). Specifically, Ouyang et al. (2022) use a pretrained language model for the reward model, opting for a smaller model relative to the policy to mitigate unstable learning. Notably, while they apply supervised fine-tuning to the policy before the RLHF phase, the reward model is initialized directly from the language model without any preliminary fine-tuning. This approach’s applicability extends beyond language models to other areas. A notable example is Abramson et al. (2022), who, in the control domain, begin by training a policy using contrastive self-supervised learning and behavioral cloning. They then add an MLP head to the policy for the prediction of cumulative rewards.
@@
A compact representation that captures all relevant information for human preferences while minimizing noise can greatly enhance preference learning efficiency. It may also generalize better than a representation learned end-to-end as part of the preference learning task, which may contain spurious correlations. Bobu et al. (2022) address this by proposing the learning of features through explicit human feedback using feature traces. Feature traces (see Section˜3.2.3) involve human labelers explicitly teaching relevant features one by one by demonstrating behavior in which the feature monotonically increases or decreases. This method directly aligns the learned representation with human-identified features, enhancing preference learning efficiency but requiring detailed human input. However, feature traces require labelers to be able to identify and articulate relevant features, which can be challenging. Bobu et al. (2023) offer an alternative approach with their Similarity-based Implicit Representation Learning (SIRL) method. SIRL learns representations from similarity queries (see Section˜3.2.3), where human labelers provide feedback on whether behaviors are similar or different concerning the features that matter to them. This method captures a broader range of human notions of similarity without needing explicit feature knowledge, thus reducing the cognitive load on human labelers. In summary, while both approaches emphasize human feedback’s centrality in representation learning, they differ in their methods of gathering this feedback. The feature traces used by Bobu et al. (2022) require specific feature knowledge, whereas SIRL used by Bobu et al. (2023) uses more intuitive similarity assessments, potentially offering a more user-friendly way to capture human preferences.
These diverse methods of using prior data demonstrate the potential for enhancing data efficiency in RLHF, enabling more effective learning from limited human feedback.